# PENEX: AdaBoost-Inspired Neural Network Regularization

### v1

@@ 2.1 Theoretical Properties
$\rho_{*}(\alpha,f)\;=\;\alpha\frac{\mathbb{E}[\mathcal{L}_{\mathrm{\scriptscriptstyle EX}}(f;\alpha)]}{\mathbb{E}[\sum_{j=1}^{K}f^{(j)}(\mathbf{x})]}.$ (7)
@@ 2.2 Practical Implementation
5 $\rho^{\prime}\leftarrow\alpha\frac{\mathcal{L}_{\mathrm{\scriptscriptstyle EX}}^{\mathcal{D}_{\mathrm{b}}}(f;\alpha)}{\hat{\mathbb{E}}_{\mathcal{D}_{\mathrm{b}}}[\sum_{j=1}^{K}f^{(j)}(\mathbf{x})]+\epsilon}$ // 7
@@ 2.3 Relationship to Multi-Class AdaBoost
We now proceed to demonstrate formally that gradient descent on PENEX is analaguous to AdaBoost in that gradient descent steps can be thought of as constrained parameters of weak learners that minimize the exponential loss. To see this, we consider a linearization about the previous parameters $\theta_{m-1}$, which assumes additivity as an approximation:
@@ 5 Discussion & Limitations
The only experiment were PENEX does not perform better or equally well in comparison to other methods is the ImageNet experiment, indicating that PENEX is advantageous only in low-data regimes, such as CIFAR-100. We believe that developing remedies for this issue is another relevant direction for future research.
@@ Appendix A Proofs
Appendix B Stategewise Additive Modeling using a Multi-class Exponential loss
@@ C.1 Inference
Motivated by the Fisher consistency result (Prop. 2.1), we always rescale logits that are optimized on PENEX by $1-\alpha$. Specifically, to obtain a prediction $\hat{P}_{\scriptscriptstyle\textrm{PENEX}}$, we always use
$\hat{P}_{\scriptscriptstyle\textrm{PENEX}}(y\,|\,\mathbf{x})\;\coloneqq\;\frac{\exp\{(1-\alpha)f^{(y)}(\mathbf{x})\}}{\sum_{j=1}^{K}\exp\{(1-\alpha)f^{(j)}(\mathbf{x})\}},$ (20)
@@ D.6.1 Computer Vision
We do not perform data augmentation for any computer vision task.

### v2

@@ 2.1 Theoretical Properties
$\rho_{*}(\alpha,f)\;=\;\alpha\frac{\mathbb{E}[\mathcal{L}_{\mathrm{\scriptscriptstyle EX}}(f;\alpha)]}{\mathbb{E}[\mathrm{SE}(f;\mathbf{x})]}.$ (7)
@@ 2.2 Practical Implementation
5 $\rho^{\prime}\leftarrow\alpha\frac{\mathcal{L}_{\mathrm{\scriptscriptstyle EX}}^{\mathcal{D}_{\mathrm{b}}}(f;\alpha)}{\hat{\mathbb{E}}_{\mathcal{D}_{\mathrm{b}}}[\mathrm{SE}(f;\mathbf{x})]+\epsilon}$ // 7
@@ 2.3 Relationship to Multi-Class AdaBoost
We now proceed to demonstrate formally that gradient descent on PENEX is analogous to AdaBoost in that gradient descent steps can be thought of as constrained parameters of weak learners that minimize the exponential loss. To see this, we consider a linearization about the previous parameters $\theta_{m-1}$, which assumes additivity as an approximation:
@@ 5 Discussion & Limitations
The only experiment where PENEX does not perform better or equally well in comparison to other methods is the ImageNet experiment, indicating that PENEX is advantageous only in low-data regimes, such as CIFAR-100. We believe that developing remedies for this issue is another relevant direction for future research.
@@ Appendix A Proofs
Appendix B Stagewise Additive Modeling using a Multi-class Exponential loss
@@ C.1 Inference
Motivated by the Fisher consistency result (Prop. 2.1), we always rescale logits that are optimized on PENEX by $1+\alpha$. Specifically, to obtain a prediction $\hat{P}_{\scriptscriptstyle\textrm{PENEX}}$, we always use
$\hat{P}_{\scriptscriptstyle\textrm{PENEX}}(y\,|\,\mathbf{x})\;\coloneqq\;\frac{\exp\{(1+\alpha)f^{(y)}(\mathbf{x})\}}{\sum_{j=1}^{K}\exp\{(1+\alpha)f^{(j)}(\mathbf{x})\}},$ (20)
@@ D.6.1 Computer Vision
We do not perform data augmentation for any computer vision task, except for ImageNet.